Evaluate each expression.....
7. 6P4

8. 5(3P2)

9. 5C2 + 5C1

10. 7C3/7C4
11. There are 14 different types of boxed nails in the hardware store. You plan to buy 5 different types of boxed nails. How many different combinations are there?


12. In how many different orders can six colored blocks be chosen from a set of 23 different blocks?

Answer by stanbon(48564)

Evaluate each expression.....
7. 6P4 = 6!/(6-4)! = 6!/2! = 6*5*4*3 = 360
-------------------------------------
8. 5(3P2) = 5[3!/(3-2)!] = 5[3!/1!] = 5*3*2*1 = 6
-------------------------------------
9. 5C2 + 5C1 = [5*4/1*2] + [5/1] = 10+5 = 15
-------------------------------------
10. 7C3/7C4 = [7*6*5/1*2*3]/[7*6*5*4/1*2*3*4] = 1
================================
11. There are 14 different types of boxed nails in the hardware store. You plan to buy 5 different types of boxed nails. How many different combinations are there?
14C5 = [14*13*12*11*10]/[1*2*3*4*5] = 2002
--------------------------------
12. In how many different orders can six colored blocks be chosen from a set of 23 different blocks?
23C6 = 100947
==================
Cheers,
Stan H.